# Teaching LLMs How to Learn with Contextual Fine-Tuning

### v1

@@ Abstract
Prompting Large Language Models (LLMs), or providing context on the expected model of operation, is an effective way to steer the outputs of such models to satisfy human desiderata after they have been trained. But in rapidly evolving domains, there is often need to fine-tune LLMs to improve either the kind of knowledge in their memory or their abilities to perform open ended reasoning in new domains. When human’s learn new concepts, we often do so by linking the new material that we are studying to concepts we have already learned before. To that end, we ask, can prompting help us teach LLMs how to learn. In this work, we study a novel generalization of instruction tuning, called contextual fine-tuning, to fine-tune LLMs. Our method leverages prompts designed to mimic human cognitive strategies in learning and problem-solving to guide the learning process during training, aiming to improve the model’s interpretation and understanding of domain-specific knowledge. We empirically demonstrate that this simple yet effective modification improves the ability of LLMs to be fine-tuned rapidly on new datasets both within the medical and financial domains. Project page: https://younwoochoi.github.io/cft-iclr/
@@ 1 Introduction
However, LLMs remain unaware of information and events occurring after their knowledge cutoff; in fast-moving domains and in scenarios where deployment requires knowledge of up-to-date information, there is a need to remedy this limitation. There are two popular approaches to this problem. The first is to increase the context length of the model until all anticipated new information fits within this context (the largest of which is Google’s Gemini-1.5 model (Reid et al., 2024) with a context length of two million tokens). However, even context lengths this large can be exhausted and it is unclear whether the model’s attention mechanism is capable of accurately inferring signal regardless of where it is in the context. The alternate approach uses external knowledge stores via retrieval augmented systems (Lewis et al., 2021). This approach works well when the reasoning abilities already learned by the model suffice to process and extract the relevant information. But gradient-based learning remains vital in scenarios where there is a need to teach the model how to manipulate new tools or learn new strategies for reasoning.
@@
The simplest approach to update model knowledge via fine-tuning is to continue pretraining the base model. Unfortunately, the data and training procedure necessary to replicate the additional fine-tuning and alignment phases are rarely open-sourced in chat-based models. The general practice is to fine-tune the aligned model with new domain-specific knowledge. Models that have undergone instruction fine-tuning and alignment training are more amenable to interacting with users but are harder to update with new knowledge. But training to update the knowledge can result in catastrophic forgetting of knowledge gained during pretraining, or loss of capabilities like instruction-following and task-solving (Wang et al., 2023a). There is a need to develop methods that enable us to quickly improve reasoning and recall in aligned LLMs for domain specific fine-tuning.
@@
    To study the impact of our method, we create two datasets in the biomedical domain: the first consisting of 121,489 journal articles from 37 diverse topics in biology and medicine, and second comprising 29 open-source medical textbooks.
@@
    We show that contextual fine-tuning can be used to update a model’s knowledge more efficiently than continued pretraining and instruction tuning. We show increased performance on both real-world dataset and Q&A tasks while using carefully constructed synthetic data to better understand where performance gains arise from.
@@ 2 Related Work
The common paradigm used in training instruction aligned (”chat”) LLMs involves three steps: pretraining on unlabeled corpora, performing instruction tuning, followed by reward-based preference training, as used in Ouyang et al. (2022). The instruction tuning phase is generally used as a first step to aligning LLMs to human instructions or when access to human-labeled preference data is limited. Instruction tuning significantly narrows the divide between models’ traditional next-word prediction objectives and the practical need for models to adhere to explicit human instructions. Wei et al. (2022a) have highlighted how this approach markedly boosts zero-shot performance across previously unseen tasks, underlining its effectiveness. Earlier work (Wei et al., 2022a; Iyer et al., 2022) have proposed using a large set of instructions for NLP tasks, while more recent findings (Wang et al., 2023c; Taori et al., 2023; Peng et al., 2023; Zhou et al., 2024) have found success using increasingly smaller and higher quality instruction datasets on open sourced pretrained models such as Llama (Touvron et al., 2023a). Notably, as Gudibande et al. (2023) discovers, training on instructions in the instruction tuning phase does not improve the underlying capabilities of the models; these models merely imitate the instruction following template. Recent developments have explored alternative approaches to standard instruction tuning. Wang et al. (2025) propose Critique Fine-Tuning, where models learn to critique noisy responses rather than imitate correct ones, demonstrating improvements over SFT across mathematical reasoning benchmarks.
The primary distinction between contextual fine-tuning and instruction fine-tuning lies in the semantic content of the tokens used as input. Instruction tuning has input-output pairs ($x,y$) that are data point specific (e.g., $x$ = ”Who is the current president of the United States?”, $y$ = ”Joe Biden”). Within instruction tuning, there is a specific, narrow question for which there exists a right answer that the model is expected to identify. In contextual fine-tuning, our intent is to pair $y$ with a randomly sampled contextual prompt $x$ which serves as guidance for the model to learn the most important information. $x$ can be specific or general desiderata useful for learning intended to prime the model to contextualize and incorporate the knowledge in $y$ within its parameters.
@@
While pretraining on trillions of unlabeled tokens creates generalist foundation models (Bommasani et al., 2021; Touvron et al., 2023a; b; Anil et al., 2023; OpenAI et al., 2024), injecting domain-specific expertise into models while retaining the generalist remains an active front of research. Improving the underlying capabilities of LLMs is a more difficult challenge than simply aligning to instructions, in part due to the much larger dataset requirement. While smaller LLMs are capable of outperforming the larger monoliths in specific domains (Gunasekar et al., 2023; Li et al., 2023), or pushing language modeling in a simplified domain to the extreme (Eldan & Li, 2023). Several lines of work have explored using better datasets to improve performance on existing pretrained models as a form of continual learning in areas such as medicine, finance, chip design, coding, and mathematics (Chen et al., 2023b; Wu et al., 2023; Liu et al., 2023; Rozière et al., 2024; Paster et al., 2023; Azerbayev et al., 2023). Such continued training runs the risk of model forgetting, in which general reasoning and capabilities in other domains tend to decline (Luo et al., 2023; Wang et al., 2023a; Chen et al., 2023a).
@@ 3.1 Notation & Background
Continued pretraining (CPT) leverages large volumes of unlabeled domain-specific data to refine the model’s understanding of the domain. Given sequences of tokens $x=(x_{1},x_{2},\ldots,x_{n})$ sampled from $\mathcal{D}_{train}^{raw}$, the model is trained using the causal language modeling objective, which predicts the next token given the previous tokens.
@@ 3.2 Contextual Fine-tuning
    • ”Concentrate on understanding the core principles and essential facts in the following text. Pay special attention to definitions, examples, and conclusions.”
@@
    • ”As you read the next passage, relate its content to its broader context and implications. Think about how this information connects to what you’ve learned previously.”
@@
    • ”Critically analyze the upcoming information. Look for underlying assumptions, evaluate arguments, and consider different perspectives.”
@@
    • ”Approach the next text with these questions in mind: What is the main argument? How is evidence used to support it? What are the implications of these findings?”
The selection of prompts inspired by educational theories satisfy two important criteria – they are compact and general-purpose (in that they are applicable across different domains). Indeed, these prompts need not necessarily be optimal, rather they need only contain semantic content effective for modifying the gradients during the process of fine-tuning.
@@ 4 OpenMedText
To evaluate the effectiveness of contextual fine-tuning in a domain-adaptive setting, we curated a dataset consisting of both academic journal articles and educational textbooks. Our objective was to assemble a corpus that not only covers a wide range of topics within bio-medicine but also provides structured textual data (of varying levels of quality) suitable to align with our goal of studying how well LLMs can learn using contextual prompts. The inclusion of textbooks provides structured and pedagogically organized content, which is conducive to the learning processes we aim to emulate.
@@
Medical Textbooks: We also curated 29 open-source medical textbooks into our dataset. Textbooks were chosen because they provide structured, comprehensive overviews of medical knowledge, organized pedagogically to facilitate learning. This aligns with our objective of leveraging contextual fine-tuning to enhance the learning processes of language models, as textbooks inherently contain explanations, definitions, and educational narratives beneficial for model training. The data we collect have the following characteristics:
@@ 5.1 Experimental Setup
The effectiveness of the fine-tuning approach in each domain is evaluated using several domain-specific benchmarks. In the financial domain, we consider (1) the sentiment analysis task FiQA (Xie et al., 2023) where LLMs predict sentiments categorized as ’positive’, ’neutral’, or ’negative’ in financial texts. (2) The headline classification task MultiFin (Jørgensen et al., 2023; Xie et al., 2023), where LLMs categorize each news article into one of six categories based on the headline. (3) Causal20 (Xie et al., 2023), which involves classifying sentences extracted from financial news as either depicting a ’causal’ or ’noise’ relationship between financial events. For the biomedical domain, we consider the following multiple-choice question (MCQ) datasets from Massive Multitask Language Understanding (MMLU) (Hendrycks et al., 2021): (1) Anatomy, (2) Clinical Knowledge, (3) College Biology, (4) College Medicine, (5) Medical Genetics, and (6) Professional Medicine. We also use MedQA, a collection of multiple-choice questions from the professional medical board exams (Jin et al., 2020).
@@
We use several configurations of the Llama-2 models to evaluate the effectiveness of contextual fine-tuning. More details can be found under Appendix A.3. In our evaluation, MCQs are formatted with questions followed by several options labeled with ID symbols (e.g., A/B/C/D). Building on the approach outlined in Zheng et al. (2024), we instruct the language models to predict an option ID symbol rather than the textual content of the answer. This method addresses a critical issue: the likelihood of the answer’s text being naturally plausible could be conflated with its likelihood of being the correct response due to the model’s linguistic biases. However Robinson & Wingate (2023) raises concerns regarding LLMs’ inherent selection biases, which highlights that these models may show a preference for specific option IDs. To counteract this bias and enhance the validity of our evaluations, we adopt one of debiasing techniques as prescribed in the aforementioned work. See Appendix C.1 for the detailed method.
@@ 5.2 Results
We evaluate the zero-shot performance of large language models (LLMs) trained with different methods across both medical and financial benchmarks. See Appendix E.3 for extended results on general-purpose and instruction-following benchmarks. It is important to note that our primary objective is not to achieve state-of-the-art performance but to assess the relative improvements offered by contextual fine-tuning (CFT), instruction fine-tuning (IFT), and continued pretraining (CPT) compared to a baseline model. We focus primarily on a metric for evaluating the effectiveness of these training approaches: %$\%\Delta^{CFT}_{CPT}$, which denotes the performance difference between CFT and CPT.
@@
Table 1: Medical Benchmarks (Zero-shot). The results show that the 7B model achieved a %$\%\Delta^{CFT}_{CPT}$ of 1.85%. The 13B model demonstrated increased effectiveness with a %$\%\Delta^{CFT}_{CPT}$ of 2.25%, indicating that CFT’s impact grows with the model’s scale.
@@
We asses how contextual fine-tuning performs across different model scales. Table 1 presents medical benchmarks for the 7B and 13B model scales which shows a %$\%\Delta^{CFT}_{CPT}=1.85\%$. For the 13B model, these metrics increase to %$\%\Delta^{CFT}_{CPT}=2.25\%$. Similarly, Tables 3 and 3 contain financial benchmarks. The 7B model records a %$\%\Delta^{CFT}_{CPT}=4.32\%$ whereas the 13B model, the results are %$\%\Delta^{CFT}_{CPT}=3\%$. The results demonstrate that the simple augmentation of contextual prompting can help increase performance across the board.
@@
The tables in Appendix D.1 show the performance on the medical and financial benchmarks while holding the model scale constant. The base non-instruct model holds an average accuracy of 41.34% on the medical benchmarks. Our experiments find that combining training schemes provides the greatest boost in fine-tuning performance. In particular, combining CFT and IFT gives a performance boost of %$\%\Delta^{CFT+IFT}_{Base}=2.95\%$ compared to %$\%\Delta^{CPT+IFT}_{Base}=1.91\%$. Similar trends are seen in the financial benchmarks where the same combination led to an increase of %$\%\Delta^{CFT+IFT}_{Base}=36.28\%$ in F1 score. These results further solidify that augmenting the CPT stage of fine-tuning to instead use CFT provides a near-free boost in performance. More detailed analyses can be found in Appendix D.1. Additionally, we compare CFT with AdaptLLM (Cheng et al., 2024), a method of domain-specific continued pretraining on medical benchmarks. Table 4 shows that CFT consistently achieves higher accuracy across all tasks by 4.89% on average. This result highlights CFT’s superior effectiveness on our full-text medical datasets compared to AdaptLLM. Refer to Appendix E.2 for more details on the AdaptLLM methodology.
The semantic content of the contextual prompts are important to improving performance.
The core aspect of our study involves examining the impact of additional context on model performance, and specifically how the signal from this context provides a boost in learning performance. We conduct an ablation by introducing negative contextual prompts, which are designed to mislead the model by suggesting that the following information is incorrect. These results can be found in tables under Appendix D.2. In the financial domain, the impact of negative prompts is evident. The 7B model experienced a performance drop of %$\%\Delta^{-CFT}_{CFT}=-3.41\%$, and the 13B model sees a decrease of %$\%\Delta^{-CFT}_{CFT}=-2.39\%$. All models undergoing negative contextual fine-tuning still perform better than those subjected to CPT. In addition to experiments with negative contextual prompts, we investigate text-adaptive Contextual Fine-Tuning (TextAdaptCFT) using automatically generated, text-dependent prompts. As shown in Table 5, TextAdaptCFT achieves an average accuracy of 46.31%, outperforming both the baseline Chat model (45.05%) and the Chat CPT model (45.96%). This suggests that our manually constructed contextual prompts are not the only viable solution, and that semantic content within prompts plays a crucial role in enhancing model performance. Refer to Appendix D.3 for additional details on our automated prompt construction process.

### v2

@@ Abstract
Prompting Large Language Models (LLMs), or providing context on the expected model of operation, is an effective way to steer the outputs of such models to satisfy human desiderata after they have been trained. But in rapidly evolving domains, there is often a need to fine-tune LLMs to improve either the kind of knowledge in their memory or their abilities to perform open-ended reasoning in new domains. When humans learn new concepts, we often do so by linking the new material that we are studying to concepts we have already learned before. To that end, we ask, can prompting help us teach LLMs how to learn? In this work, we study a novel generalization of instruction tuning, called contextual fine-tuning, to fine-tune LLMs. Our method leverages prompts designed to mimic human cognitive strategies in learning and problem-solving to guide the learning process during training, aiming to improve the model’s interpretation and understanding of domain-specific knowledge. We empirically demonstrate that this simple yet effective modification improves the ability of LLMs to be fine-tuned rapidly on new datasets both within the medical and financial domains. Project page: https://younwoochoi.github.io/cft-iclr/
@@ 1 Introduction
However, LLMs remain unaware of information and events occurring after their knowledge cutoff; in fast-moving domains and in scenarios where deployment requires knowledge of up-to-date information, there is a need to remedy this limitation. There are two popular approaches to this problem. The first is to increase the context length of the model until all anticipated new information fits within this context (the largest of which is Google’s Gemini-1.5 model (Reid et al., 2024) with a context length of two million tokens). However, even context lengths this large can be exhausted, and it is unclear whether the model’s attention mechanism is capable of accurately inferring the signal regardless of where it is in the context. The alternative approach utilizes external knowledge stores via retrieval-augmented systems (Lewis et al., 2021). This approach works well when the reasoning abilities already learned by the model suffice to process and extract the relevant information. Gradient-based learning, however, remains vital in scenarios where there is a need to teach the model how to manipulate new tools or learn new strategies for reasoning.
@@
The simplest approach to update model knowledge via fine-tuning is to continue pretraining the base model. Unfortunately, the data and training procedure necessary to replicate the additional fine-tuning and alignment phases are rarely open-sourced in chat-based models. The general practice is to fine-tune the aligned model with new domain-specific knowledge. Models that have undergone instruction fine-tuning and alignment training are more amenable to interacting with users but are harder to update with new knowledge. However, training to update knowledge can result in catastrophic forgetting of knowledge gained during pretraining, or a loss of capabilities such as instruction-following and task-solving (Wang et al., 2023a). There is a need to develop methods that enable us to quickly improve reasoning and recall in aligned LLMs for domain specific fine-tuning.
@@
    To study the impact of our method, we create two datasets in the biomedical domain: the first consisting of 121,489 journal articles from 37 diverse topics in biology and medicine, and the second comprising 29 open-source medical textbooks.
@@
    We show that contextual fine-tuning can be used to update a model’s knowledge more efficiently than continued pretraining and instruction tuning. We show increased performance on both real-world datasets and Q&A tasks while using carefully constructed synthetic data to better understand where performance gains arise from.
@@ 2 Related Work
The common paradigm used in training instruction-aligned (“chat”) LLMs involves three steps: pretraining on unlabeled corpora, performing instruction tuning, followed by reward-based preference training, as used in Ouyang et al. (2022). The instruction tuning phase is generally used as a first step to aligning LLMs to human instructions or when access to human-labelled preference data is limited. Instruction tuning significantly narrows the divide between models’ traditional next-word prediction objectives and the practical need for models to adhere to explicit human instructions. Wei et al. (2022a) have highlighted how this approach markedly boosts zero-shot performance across previously unseen tasks, underlining its effectiveness. Earlier work (Wei et al., 2022a; Iyer et al., 2022) have proposed using a large set of instructions for NLP tasks, while more recent findings (Wang et al., 2023c; Taori et al., 2023; Peng et al., 2023; Zhou et al., 2024) have found success using increasingly smaller and higher quality instruction datasets on open sourced pretrained models such as Llama (Touvron et al., 2023a). Notably, as Gudibande et al. (2023) discovers, training on instructions in the instruction tuning phase does not improve the underlying capabilities of the models; these models merely imitate the instruction following template. Recent developments have explored alternative approaches to standard instruction tuning. Wang et al. (2025) propose Critique Fine-Tuning, where models learn to critique noisy responses rather than imitate correct ones, demonstrating improvements over SFT across mathematical reasoning benchmarks.
The primary distinction between contextual fine-tuning and instruction fine-tuning lies in the semantic content of the tokens used as input. Instruction tuning has input-output pairs ($x,y$) that are data point specific (e.g., $x$ = “Who is the current president of the United States?”, $y$ = “Joe Biden”). Within instruction tuning, there is a specific, narrow question for which there exists a right answer that the model is expected to identify. In contextual fine-tuning, our intent is to pair $y$ with a randomly sampled contextual prompt $x$, which serves as guidance for the model to learn the most important information. $x$ can be specific or general desiderata useful for learning, intended to prime the model to contextualize and incorporate the knowledge in $y$ within its parameters.
@@
While pretraining on trillions of unlabeled tokens creates generalist foundation models (Bommasani et al., 2021; Touvron et al., 2023a; b; Anil et al., 2023; OpenAI et al., 2024), injecting domain-specific expertise into models while retaining the generalist remains an active front of research. Improving the underlying capabilities of LLMs is a more difficult challenge than simply aligning to instructions, in part due to the much larger dataset requirement. While smaller LLMs are capable of outperforming the larger monoliths in specific domains (Gunasekar et al., 2023; Li et al., 2023), or pushing language modelling in a simplified domain to the extreme (Eldan & Li, 2023). Several lines of work have explored using better datasets to improve performance on existing pretrained models as a form of continual learning in areas such as medicine, finance, chip design, coding, and mathematics (Chen et al., 2023b; Wu et al., 2023; Liu et al., 2023; Rozière et al., 2024; Paster et al., 2023; Azerbayev et al., 2023). Such continued training runs the risk of model forgetting, in which general reasoning and capabilities in other domains tend to decline (Luo et al., 2023; Wang et al., 2023a; Chen et al., 2023a).
@@ 3.1 Notation & Background
Continued pretraining (CPT) leverages large volumes of unlabeled domain-specific data to refine the model’s understanding of the domain. Given sequences of tokens $x=(x_{1},x_{2},\ldots,x_{n})$ sampled from $\mathcal{D}_{train}^{raw}$, the model is trained using the causal language modelling objective, which predicts the next token given the previous tokens.
@@ 3.2 Contextual Fine-tuning
    • “Concentrate on understanding the core principles and essential facts in the following text. Pay special attention to definitions, examples, and conclusions.”
@@
    • “As you read the next passage, relate its content to its broader context and implications. Think about how this information connects to what you’ve learned previously.”
@@
    • “Critically analyze the upcoming information. Look for underlying assumptions, evaluate arguments, and consider different perspectives.”
@@
    • “Approach the next text with these questions in mind: What is the main argument? How is evidence used to support it? What are the implications of these findings?”
The selection of prompts inspired by educational theories satisfies two important criteria – they are compact and general-purpose (in that they are applicable across different domains). Indeed, these prompts need not necessarily be optimal, rather they need only contain semantic content effective for modifying the gradients during the process of fine-tuning.
@@ 4 OpenMedText
To evaluate the effectiveness of contextual fine-tuning in a domain-adaptive setting, we curated a dataset consisting of both academic journal articles and educational textbooks. Our objective was to assemble a corpus that not only covers a wide range of topics within biomedicine but also provides structured textual data (of varying levels of quality) suitable for aligning with our goal of studying how well LLMs can learn using contextual prompts. The inclusion of textbooks provides structured and pedagogically organized content, which is conducive to the learning processes we aim to emulate.
@@
Medical Textbooks: We also curated 29 open-source medical textbooks into our dataset. Textbooks were chosen because they provide structured, comprehensive overviews of medical knowledge, organized pedagogically to facilitate learning. This aligns with our objective of leveraging contextual fine-tuning to enhance the learning processes of language models, as textbooks inherently contain explanations, definitions, and educational narratives beneficial for model training. The data we collect has the following characteristics:
@@ 5.1 Experimental Setup
The effectiveness of the fine-tuning approach in each domain is evaluated using several domain-specific benchmarks. In the financial domain, we consider (1) the sentiment analysis task FiQA (Xie et al., 2023) where LLMs predict sentiments categorized as ‘positive’, ‘neutral’, or ‘negative’ in financial texts. (2) The headline classification task MultiFin (Jørgensen et al., 2023; Xie et al., 2023), where LLMs categorize each news article into one of six categories based on the headline. (3) Causal20 (Xie et al., 2023), which involves classifying sentences extracted from financial news as either depicting a ‘causal’ or ‘noise’ relationship between financial events. For the biomedical domain, we consider the following multiple-choice question (MCQ) datasets from Massive Multitask Language Understanding (MMLU) (Hendrycks et al., 2021): (1) Anatomy, (2) Clinical Knowledge, (3) College Biology, (4) College Medicine, (5) Medical Genetics, and (6) Professional Medicine. We also use MedQA, a collection of multiple-choice questions from the professional medical board exams (Jin et al., 2020).
@@
We use several configurations of the Llama-2 models to evaluate the effectiveness of contextual fine-tuning. More details can be found under Appendix A.3. In our evaluation, MCQs are formatted with questions followed by several options labelled with ID symbols (e.g., A/B/C/D). Building on the approach outlined in Zheng et al. (2024), we instruct the language models to predict an option ID symbol rather than the textual content of the answer. This method addresses a critical issue: the likelihood of the answer’s text being naturally plausible could be conflated with its likelihood of being the correct response due to the model’s linguistic biases. However, Robinson & Wingate (2023) raises concerns regarding LLMs’ inherent selection biases, which highlights that these models may show a preference for specific option IDs. To counteract this bias and enhance the validity of our evaluations, we adopt one of the debiasing techniques as prescribed in the aforementioned work. See Appendix C.1 for the detailed method.
@@ 5.2 Results
We evaluate the zero-shot performance of large language models (LLMs) trained with different methods across both medical and financial benchmarks. See Appendix E.3 for extended results on general-purpose and instruction-following benchmarks. It is important to note that our primary objective is not to achieve state-of-the-art performance but to assess the relative improvements offered by contextual fine-tuning (CFT), instruction fine-tuning (IFT), and continued pretraining (CPT) compared to a baseline model. We focus primarily on a metric for evaluating the effectiveness of these training approaches: $\%\Delta^{CFT}_{CPT}$, which denotes the performance difference between CFT and CPT.
@@
Table 1: Medical Benchmarks (Zero-shot). The results show that the 7B model achieved a $\%\Delta^{CFT}_{CPT}$ of 1.85%. The 13B model demonstrated increased effectiveness with a $\%\Delta^{CFT}_{CPT}$ of 2.25%, indicating that CFT’s impact grows with the model’s scale.
@@
We asses how contextual fine-tuning performs across different model scales. Table 1 presents medical benchmarks for the 7B and 13B model scales, which show a $\%\Delta^{CFT}_{CPT}=1.85\%$. For the 13B model, these metrics increase to $\%\Delta^{CFT}_{CPT}=2.25\%$. Similarly, Tables 3 and 3 contain financial benchmarks. The 7B model records a $\%\Delta^{CFT}_{CPT}=4.32\%$ whereas the 13B model, the results are $\%\Delta^{CFT}_{CPT}=3\%$. The results demonstrate that the simple augmentation of contextual prompting can help increase performance across the board.
@@
The tables in Appendix D.1 show the performance on the medical and financial benchmarks while holding the model scale constant. The base non-instruct model holds an average accuracy of 41.34% on the medical benchmarks. Our experiments show that combining training schemes yields the greatest improvement in fine-tuning performance. In particular, combining CFT and IFT gives a performance boost of $\%\Delta^{CFT+IFT}_{Base}=2.95\%$ compared to $\%\Delta^{CPT+IFT}_{Base}=1.91\%$. Similar trends are seen in the financial benchmarks where the same combination led to an increase of $\%\Delta^{CFT+IFT}_{Base}=36.28\%$ in F1 score. These results further solidify that augmenting the CPT stage of fine-tuning to instead use CFT provides a near-free boost in performance. More detailed analyses can be found in Appendix D.1. Additionally, we compare CFT with AdaptLLM (Cheng et al., 2024), a method of domain-specific continued pretraining on medical benchmarks. Table 4 shows that CFT consistently achieves higher accuracy across all tasks by 4.89% on average. This result highlights CFT’s superior effectiveness on our full-text medical datasets compared to AdaptLLM. Refer to Appendix E.2 for more details on the AdaptLLM methodology.
The semantic content of the contextual prompts is important to improving performance.
The core aspect of our study involves examining the impact of additional context on model performance, specifically how the signal from this context enhances learning performance. We conduct an ablation by introducing negative contextual prompts, which are designed to mislead the model by suggesting that the following information is incorrect. These results are presented in tables in Appendix D.2. In the financial domain, the impact of negative prompts is evident. The 7B model experienced a performance drop of $\%\Delta^{-CFT}_{CFT}=-3.41\%$, and the 13B model sees a decrease of $\%\Delta^{-CFT}_{CFT}=-2.39\%$. All models undergoing negative contextual fine-tuning still outperform those subjected to CPT. In addition to experiments with negative contextual prompts, we investigate text-adaptive Contextual Fine-Tuning (TextAdaptCFT) using automatically generated, text-dependent prompts. As shown in Table 5, TextAdaptCFT achieves an average accuracy of 46.31%, outperforming both the baseline Chat model (45.05%) and the Chat CPT model (45.96%). This suggests that our manually constructed contextual prompts are not the only viable solution, and that semantic content within prompts plays a crucial role in enhancing model performance. Refer to Appendix D.3 for additional details on our automated prompt construction process.